OpenStudy (anonymous):

Find all the real square roots of -9/16.

OpenStudy (helder_edwin):

negative numbers do not have real square roots

OpenStudy (solomonzelman):

real sqrt? well the square rot of this is \[±i3/4\]

OpenStudy (anonymous):

no real root -3/4 -3/4 and 3/4 81/256

OpenStudy (solomonzelman):

It is an imaginary sqrt though.

OpenStudy (solomonzelman):

These choices are INCORRECT!

OpenStudy (solomonzelman):

\[\huge {\sqrt{-\frac{9}{16}} }=±i\frac{3}{4}\]

OpenStudy (solomonzelman):

 not -9/16